### v2

@@ Abstract
Transmission line defect detection presents persistent challenges for automated inspection systems due to the prevalence of small-scale defects, complex backgrounds, and varying illumination conditions in UAV-captured imagery. While existing deep learning methods have achieved considerable progress using RGB images alone, they remain fundamentally limited in distinguishing geometrically subtle defects from visually similar background elements, particularly when chromatic contrast is insufficient. This paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network that integrates RGB appearance and depth geometry through a principled “purify-then-fuse” paradigm for robust defect detection. The architecture comprises two complementary mechanisms: a Semantic Recomposition Module that performs dictionary-based feature purification by projecting representations into a learned codebook with controllable capacity, thereby filtering modality-specific noise while preserving defect-relevant patterns; and a Contextual Semantic Integration Framework that establishes global spatial dependencies through partial-channel attention, enabling the network to leverage structural priors for semantic disambiguation. Position-wise normalization within the purification stage transforms cross-modal alignment from implicit weight-based learning into explicit reconstruction-driven enhancement, ensuring that heterogeneous RGB and depth features attain compatible statistical distributions prior to fusion. Extensive experiments on TL-RGBD, a large-scale benchmark where 94.5% of annotated instances qualify as small objects, demonstrate that CMAFNet achieves 32.2% mAP50 and 12.5% APs, surpassing the best baseline by 9.8 and 4.0 absolute percentage points, respectively. The lightweight configuration attains 24.8% mAP50 at 228 FPS with only 4.9M parameters, outperforming all YOLO variants while matching larger transformer-based detectors at a fraction of their computational cost.
@@ 1 Introduction
Electrical power transmission networks constitute the critical infrastructure of modern society, serving as the backbone for energy delivery across industrial, commercial, and residential sectors (Zhang et al., 2024b; Luo et al., 2023). Transmission lines, as the most exposed and geographically dispersed components of power grids, are continuously subjected to harsh environmental conditions, mechanical stress, and material degradation over extended operational periods (Shen et al., 2025; Deng et al., 2024). Common defects—including insulator contamination, conductor corrosion, fitting damage, and bird nest interference—pose substantial risks to grid reliability and public safety if left undetected (Bao et al., 2022; Zhang et al., 2025a; Cheng and Liu, 2024). The timely and accurate identification of such defects is therefore essential not only for preventing catastrophic failures and large-scale blackouts but also for optimizing maintenance scheduling and extending asset lifespan (Yang et al., 2023). Given the vast scale of transmission networks and the increasing demand for grid resilience, automated and intelligent defect detection has emerged as a pressing research priority in both the power engineering and computer vision communities (Li et al., 2023; Zhang et al., 2025b).
Conventional approaches to transmission line defect detection have long confronted significant challenges in achieving reliable and scalable performance (Liu et al., 2020; Song et al., 2014). Manual inspection, which dominated early practices, typically relies on human operators either physically traversing transmission corridors or remotely controlling inspection platforms along predefined paths—a process that is inherently hazardous to crew safety, operationally inefficient, and cost-prohibitive for large-scale grid networks (Pouliot and Montambault, 2012; Xu et al., 2022a). These practical constraints have driven the power industry toward automated inspection systems to ensure operational stability and safety, with intelligent visual inspection emerging as a particularly promising paradigm that enables efficient defect identification while substantially reducing manual annotation burdens and field deployment risks (Wang et al., 2025; Yi et al., 2025).
The rapid proliferation of unmanned aerial vehicles equipped with high-resolution RGB cameras has substantially transformed data acquisition paradigms, enabling efficient large-scale image collection across extensive transmission corridors (Xu et al., 2022b; da Silva et al., 2020). Building upon this advancement, numerous deep learning-based detection frameworks have been proposed, predominantly leveraging convolutional neural networks such as YOLO variants with attention mechanisms for enhanced feature extraction (Wang et al., 2024), Faster R-CNN with aerial image augmentation strategies (Cao et al., 2023), and lightweight architectures incorporating multiscale feature fusion for small defect detection (Ma et al., 2025; Wu et al., 2025). Self-supervised learning methods utilizing vision transformers have further been explored to address the challenge of limited labeled defect samples in practical inspection scenarios (Zhang et al., 2024a). Despite these methodological advances, RGB-based approaches remain fundamentally constrained by their reliance on single-modal visual information, struggling with defects exhibiting low chromatic contrast against cluttered backgrounds, geometrically ambiguous damage patterns, and frequent occlusions caused by vegetation or atmospheric interference (He et al., 2025; Chen et al., 2025). A critical trade-off persists between detection accuracy and computational efficiency: high-precision models typically demand substantial computational resources measured in tens of GFLOPs, rendering them impractical for real-time deployment on resource-constrained UAV platforms, whereas lightweight alternatives often sacrifice detection performance to achieve acceptable inference speeds (Tan et al., 2020; Chen et al., 2024). Fundamentally, RGB imagery alone cannot reliably capture critical depth-related characteristics—including surface deformation, structural protrusion, and spatial displacement—that are essential for distinguishing subtle defects from benign surface variations under complex environmental conditions (Li et al., 2025; Gupta et al., 2014).
The inherent limitations of single-modal approaches have motivated the exploration of multi-modal fusion strategies that integrate RGB appearance with depth geometry to achieve more robust defect detection. According to a recent comprehensive survey on deep multi-modal data fusion (Zhao et al., 2024), existing fusion paradigms can be broadly categorized into raw-data-level fusion, hierarchical feature fusion, and decision-level fusion. Raw-data-level fusion concatenates RGB and depth channels along the input dimension before feeding them into a single encoder, which maximally preserves original information from each modality while minimizing computational cost through a unified backbone (Zhao et al., 2024). However, this approach forces the network to learn cross-modal correspondences implicitly from the earliest layers, which proves suboptimal when RGB and depth modalities exhibit fundamentally different statistical properties and semantic structures. At the opposite extreme, decision-level fusion delays cross-modal interaction until the final prediction stage, offering straightforward implementation but providing limited interpretability of multi-modal interactions and suffering from vulnerability to single-modality failures (Zhao et al., 2024). Between these two paradigms, hierarchical feature fusion has emerged as a principled framework that leverages the multi-level representations extracted from deep networks to establish cross-modal correspondences at varying levels of abstraction. By employing dedicated encoder branches for each modality and progressively aggregating their features at multiple pyramid levels, hierarchical fusion offers two key advantages: the flexibility to determine where and how many hierarchical features are fused for specific tasks, and the natural compatibility with attention mechanisms that can upgrade the connections between multi-modal features at corresponding levels (Zhao et al., 2024). Nevertheless, existing hierarchical fusion methods predominantly adopt static fusion strategies—such as fixed element-wise summation or simple concatenation followed by convolution—that lack adaptability to the dynamic variations inherent in transmission line inspection scenarios. When RGB features suffer from illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these modality-specific corruptions into the shared representation, ultimately degrading rather than enhancing detection performance.
To address the foregoing limitations, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network that integrates RGB appearance and depth geometry through a “purify-then-fuse” paradigm for transmission line defect detection. The proposed architecture proceeds in three stages: intra-modal semantic purification that suppresses noise and stabilizes feature distributions within each modality branch, cross-modal fusion with explicit semantic enhancement that constructs shared representations emphasizing common discriminative patterns while attenuating modality-private residuals, and refined semantic guidance that propagates high-quality shared semantics reliably to small-object-sensitive detection layers. This three-stage design embodies a core principle: cross-modal fusion should not operate directly on raw backbone features containing substantial modality-specific noise; instead, each modality stream must first undergo semantic purification to create favorable conditions for subsequent alignment. Two novel modules—the Semantic Dictionary Injection Refiner and the Global Semantic Refinement Framework—are strategically deployed throughout the network to realize this paradigm. Extensive experiments on TL-RGBD, a large-scale benchmark where 94.5% of annotated instances are small objects, demonstrate that CMAFNet achieves state-of-the-art performance across multiple evaluation criteria while maintaining competitive inference efficiency suitable for practical deployment.
The main contributions of this paper are summarized as follows:
@@
    We propose CMAFNet, a Cross-Modal Alignment and Fusion Network that addresses RGB-D transmission-line defect detection through a principled “purify-then-fuse” paradigm. Unlike conventional fusion strategies that operate directly on noisy backbone features, this architecture first suppresses modality-specific artifacts within each branch before constructing cross-modal representations, thereby establishing favorable statistical conditions for subsequent alignment. Extensive experiments on a challenging benchmark dominated by small objects demonstrate that CMAFNet achieves state-of-the-art performance across multiple evaluation metrics while maintaining competitive inference efficiency.
@@
    We introduce the Semantic Recomposition Module, a dictionary-based feature purification mechanism that projects representations into a learned codebook with controllable capacity. By creating an information bottleneck through structured retrieval and residual injection, SRM filters modality-specific noise—including illumination artifacts in RGB imagery and quantization errors in depth maps—while preserving discriminative patterns essential for defect recognition. The position-wise normalization embedded within SRM transforms cross-modal alignment from implicit weight-based learning into explicit reconstruction-driven enhancement, ensuring that heterogeneous RGB and depth features attain compatible statistical distributions prior to fusion.
@@
    We design the Contextual Semantic Integration Framework, a global attention mechanism that captures long-range dependencies through partial-channel multi-head self-attention. CSIF addresses the limitation that purely convolutional fusion operates within bounded receptive fields and consequently fails to exploit structural priors—such as the regular arrangement of insulator strings and spatial relationships among transmission components—for semantic disambiguation. The partial-channel design, combined with the Adaptive Semantic Regulation Mechanism, prevents the detail erosion that typically accompanies full-channel global attention, thereby preserving the fine-grained discrimination essential for small-object detection.
The remainder of this paper is organized as follows. Section 2 elaborates the proposed CMAFNet architecture, detailing the dual-branch backbone, hierarchical fusion strategy, and the design principles underlying SRM and CSIF. Section 3 describes the experimental setup, presents quantitative comparisons with state-of-the-art methods, and provides ablation analyses that validate the contribution of each architectural component. Section 4 concludes the paper with a discussion of limitations and directions for future research.
@@ 2.1 Overview of Network Architecture
The proposed Cross-Modal Alignment and Fusion Network addresses the fundamental challenges inherent in RGB-D transmission-line defect detection, where targets of interest are predominantly small and exhibit subtle visual signatures across heterogeneous modality representations. Our architecture departs from conventional single-stream detectors by introducing a dual-branch encoding paradigm coupled with progressive cross-modal fusion, thereby enabling the network to preserve modality-specific discriminative features while constructing semantically aligned multimodal representations. The overall architecture comprises three principal components: a dual parallel backbone for independent modality encoding, hierarchical fusion layers for cross-modal feature integration, and a Feature Pyramid Network augmented with Path Aggregation Network for multi-scale detection. Two specialized modules—the Semantic Recomposition Module and the Contextual Semantic Integration Framework—are strategically embedded throughout the network to address the critical issues of intra-modal feature purification and cross-modal semantic enhancement, respectively. Fig. 1 illustrates the complete network topology and the information flow across all architectural components.
@@
Figure 1: Overall architecture of the proposed Cross-Modal Alignment and Fusion Network. The dual-branch encoder independently processes RGB and depth modalities through symmetric pathways, with Semantic Recomposition Modules deployed at $P_{3}$ and $P_{4}$ scales for intra-modal semantic purification and cross-modal distribution alignment. Cross-modal fusion is performed at $P_{4}$ and $P_{5}$ levels, where the Contextual Semantic Integration Framework provides global attention-based context modeling at the deepest semantic scale. The detection head implements bidirectional feature propagation through FPN and PAN pathways for multi-scale defect detection.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone architecture employs a symmetric dual-branch design that processes RGB imagery and depth maps through independent encoding pathways. This parallel configuration is motivated by the observation that RGB and depth modalities encode fundamentally distinct visual properties: RGB channels capture texture, color gradients, and photometric variations, whereas depth maps provide geometric structure, spatial discontinuities, and distance-based saliency. Forcing these heterogeneous signals through a shared encoder would inevitably compromise the integrity of modality-specific representations, as the network would be compelled to learn entangled features that inadequately capture the unique discriminative characteristics of each input stream.

### v3

@@ Abstract
Automated defect detection in transmission lines is essential for maintaining power grid reliability, yet it remains challenging owing to the extreme prevalence of small-scale targets in UAV-acquired imagery. Existing methods rely predominantly on RGB inputs and thus depend on photometric appearance as the primary discriminative cue—a strategy that falters when defects exhibit low chromatic contrast, geometrically ambiguous damage patterns, or partial occlusion by vegetation. Depth imagery offers complementary geometric evidence, but fusing heterogeneous modalities is non-trivial: RGB and depth features differ substantially in noise characteristics and distributional properties, and naïve integration risks propagating modality-specific artifacts into the joint representation. This paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network that follows a principled purify-then-fuse paradigm for RGB-D defect detection. Within each modality branch, a Semantic Recomposition Module projects features through a learned bottleneck with position-wise normalization, suppressing sensor-specific noise—illumination artifacts in RGB and quantization errors in depth—while narrowing the inter-modal distribution gap prior to integration. At the deepest fusion level, a Contextual Semantic Integration Framework establishes global spatial dependencies through partial-channel attention, enabling the network to exploit structural priors such as the regular arrangement of insulator strings for disambiguating small defects from visually similar background elements, without the detail erosion that full-channel attention would impose. Extensive experiments on the TL-RGBD benchmark, where 94.5% of annotated instances qualify as small objects, demonstrate that CMAFNet consistently outperforms state-of-the-art detectors across all evaluation metrics. The full-scale variant achieves 32.2% mAP50 and 12.5% small-object AP, surpassing the strongest baseline by 9.8 and 4.0 percentage points respectively, while the lightweight configuration attains 24.8% mAP50 at 228 FPS with only 4.9M parameters. Ablation studies further reveal a synergistic interaction between the two modules, yielding a combined gain of 13.7% that substantially exceeds their individual contributions, and modality ablation confirms that depth provides complementary geometric cues—particularly for boundary disambiguation and low-contrast defect localization—that RGB appearance alone cannot capture.
@@ 1 Introduction
Transmission lines are among the most spatially distributed and environmentally exposed assets in modern power grids, continuously subjected to weathering, cyclic mechanical loading, and progressive material degradation such as corrosion and insulation aging (Zhang et al., 2024b; Luo et al., 2023). Defects arising from these processes—including insulator contamination, conductor corrosion, and fitting damage—as well as external hazards such as bird nesting, can lead to flashovers, line tripping, and unplanned outages, thereby increasing maintenance costs and compromising grid reliability (Bao et al., 2022; Cheng and Liu, 2024). Timely identification of such defects is therefore critical for prioritizing maintenance scheduling, reducing unnecessary patrols, and mitigating safety risks (Yang et al., 2023). Because transmission networks often extend over thousands of kilometers, this task calls for automated visual inspection methods that can operate at high frequency and low marginal cost—requirements that are difficult to meet through manual or semi-automated workflows owing to limited workforce availability, access constraints in remote corridors, and the sheer volume of infrastructure to be covered.
Conventional inspection relies on human operators traversing transmission corridors on foot or via helicopter, or deploying robotic platforms such as line-crawling robots along predefined routes (Pouliot and Montambault, 2012; Xu et al., 2022a). These approaches are hazardous, labor-intensive, and costly to sustain at the scale of a national grid. The widespread adoption of unmanned aerial vehicles (UAVs) equipped with high-resolution cameras has substantially improved data acquisition, enabling efficient imagery collection over long corridors (Xu et al., 2022b; da Silva et al., 2020) and motivating a shift toward learning-based defect detection. The imagery obtained in this manner, however, poses detection challenges that differ markedly from those encountered in general-purpose benchmarks such as COCO or VisDrone. Transmission-line scenes are dominated by repetitive structural patterns and geometrically elongated components whose dense edge responses and strong background textures severely compress the effective pixel area available for defect discrimination. Defects typically occupy only a small fraction of the image, and this difficulty is amplified by oblique viewing angles, varying shooting distances, and persistent background clutter from vegetation, sky, and metallic infrastructure. On the TL-RGBD benchmark adopted in this study, approximately 94.5% of annotated defect instances fall below the COCO small-object area threshold of $32\times 32$ pixels at the $640\times 640$ training input resolution, indicating that detection methods not explicitly designed for small objects tend to underperform in this domain.
Deep learning has driven substantial progress in transmission-line defect detection by exploiting the growing volume of UAV-acquired data. Existing approaches rely almost exclusively on RGB imagery as the sole input modality. Representative works span YOLO variants with attention-augmented backbones (Wang et al., 2024), Faster R-CNN enhanced with aerial-image augmentation strategies (Cao et al., 2023), lightweight architectures with multiscale feature fusion (Ma et al., 2025; Wu et al., 2025), and self-supervised vision transformer pipelines (Zhang et al., 2024a). A shared limitation of these methods is their dependence on photometric appearance—texture, color, and shading—as the primary discriminative cue. This dependence becomes a major bottleneck when defects exhibit low chromatic contrast against cluttered backgrounds, when damage patterns are geometrically ambiguous, or when targets are partially occluded by vegetation (He et al., 2025; Chen et al., 2025). Because RGB images encode geometry only implicitly through shading and perspective, rather than providing explicit three-dimensional surface cues, separating subtle structural deformations from benign surface variations is particularly challenging under such conditions (Li et al., 2025; Gupta et al., 2014). An additional accuracy–efficiency tension constrains practical deployment: high-capacity models impose computational costs that exceed the processing budget of typical UAV-mounted hardware at the frame rates required for real-time inspection, whereas lightweight alternatives tend to sacrifice detection sensitivity on the small targets that dominate real-world transmission-line imagery (Tan et al., 2020; Chen et al., 2024).
Integrating depth information with RGB appearance offers a promising path toward more discriminative representations. Depth maps encode relative surface geometry and spatial extent independently of illumination and texture, thereby providing complementary evidence where photometric cues alone are insufficient—for instance, by revealing surface discontinuities at damaged fittings or volumetric protrusions caused by bird nests. However, fusing heterogeneous modalities is not straightforward. RGB and depth features originate from fundamentally different sensing processes and exhibit distinct noise characteristics and distributional properties: depth maps may contain holes, quantization artifacts, and edge bleeding, while RGB images are susceptible to specular highlights and illumination variation. Direct concatenation or element-wise combination risks propagating these modality-specific artifacts into the joint representation, ultimately degrading rather than improving detection accuracy (Zhao et al., 2024). How to align and fuse such heterogeneous representations—particularly under the combined pressures of modality-specific noise and the extreme small-object prevalence that characterizes transmission-line inspection—remains insufficiently addressed for this domain.
To address these challenges, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network built upon a purify-then-fuse paradigm for RGB-D feature integration. The design is motivated by the observation that multimodal fusion operating on noisy, statistically misaligned representations tends to inherit and amplify their deficiencies rather than compensate for them. CMAFNet therefore interposes a dedicated purification stage before fusion: within each modality branch, a Semantic Recomposition Module (SRM) reconstructs input features through a structured bottleneck that suppresses modality-specific artifacts—such as depth holes and edge bleeding, or RGB specular noise—and reduces distributional mismatch between the two feature streams prior to integration. Fusion then operates on these purified representations, and a Contextual Semantic Integration Framework (CSIF) captures long-range structural dependencies that convolutional operators model only at disproportionate cost—a capability important for reasoning about spatial relationships among small defects embedded in densely repetitive structural patterns. To mitigate the detail erosion commonly associated with global attention, CSIF adopts a partial-channel design that balances global context modeling with the fine-grained spatial discrimination required for accurate small-object localization. The resulting architecture is instantiated as a scalable model family, ranging from compact configurations suitable for real-time UAV inference to full-scale variants for server-side post-processing. Experiments on the TL-RGBD benchmark show that CMAFNet consistently outperforms compared methods across evaluation metrics and model scales; for example, the full-scale variant achieves a gain of +2.8 mAP50 and +3.6 mAPS over the strongest baseline under an identical training protocol, while the lightweight variant maintains competitive accuracy at real-time throughput.
The main contributions of this paper are as follows:
@@
    We introduce the purify-then-fuse paradigm for RGB-D feature integration in small-object-dominated detection tasks. This paradigm is instantiated as CMAFNet, a scalable architecture family spanning lightweight to full-scale configurations, which achieves consistent improvements across all model scales on TL-RGBD. Systematic fusion-strategy ablations confirm that purification before fusion yields measurable gains over direct-fusion alternatives.
@@
    We propose the Semantic Recomposition Module (SRM), which suppresses modality-specific noise—including depth quantization artifacts and RGB illumination variations—and reduces cross-modal distributional mismatch through a structured reconstruction bottleneck. Applied within each modality branch and at fused feature levels, SRM brings heterogeneous RGB and depth representations into more compatible statistical properties, as validated by modality ablations and noise-sensitivity analyses presented in Section 3.
@@
    We design the Contextual Semantic Integration Framework (CSIF), a partial-channel global attention mechanism that captures long-range structural dependencies while preserving the spatial detail essential for small-object detection. By restricting global attention to a subset of feature channels and retaining the remainder for local discrimination, CSIF addresses the tension between contextual coverage and spatial precision that limits conventional full-channel attention, as demonstrated through component-level ablation.
The remainder of this paper is organized as follows. Section 2 details the proposed CMAFNet architecture, including the dual-branch backbone, hierarchical fusion strategy, and the design of SRM and CSIF. Section 3 presents the experimental setup, quantitative comparisons, and ablation analyses. Section 4 discusses limitations and future directions.
@@ 2.1 Overview of Network Architecture
Although integrating depth geometry with RGB appearance promises to strengthen defect detection, realizing this potential hinges on how the two modalities interact during feature extraction and fusion. Naively combining them—whether by early concatenation or late decision fusion—yields limited gains, because RGB and depth signals differ substantially in statistical properties and noise characteristics (Zhao et al., 2024). Early concatenation forces the network to reconcile heterogeneous distributions from the first layer onward, while late fusion defers cross-modal interaction to the prediction stage and affords little opportunity for mutual enhancement. A more principled strategy is to fuse modality-specific features at intermediate representation levels, where sufficient semantic abstraction exists to tolerate noise yet spatial resolution remains adequate for small-object localization. Existing approaches that operate at this level, however, typically employ static operations such as element-wise summation or channel concatenation followed by convolution. Such fixed fusion schemes cannot adapt to the dynamic variations inherent in aerial inspection: when RGB features are corrupted by illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these modality-specific corruptions into the shared representation, degrading detection performance rather than enhancing it.
These considerations motivate a fusion paradigm in which each modality stream first undergoes explicit noise suppression and distribution normalization before cross-modal integration, and in which the fused representation is subsequently refined through context-aware mechanisms that extend beyond the local receptive field of convolutional operations. The proposed Cross-Modal Alignment and Fusion Network (CMAFNet) instantiates this paradigm by extending the YOLO 11(Jocher et al., 2023) single-stream detector into a dual-branch architecture for RGB-D input, retaining all standard components—C3k2, SPPF, the FPN-PAN neck, and the anchor-free decoupled detection head—without modification. Three design elements distinguish CMAFNet from the baseline: (i) a Semantic Recomposition Module (SRM) that performs bottleneck-based feature purification with position-wise distribution normalization, (ii) a Contextual Semantic Integration Framework (CSIF) that introduces attention-driven global context modeling at the deepest fusion level, and (iii) a selective fusion strategy that restricts cross-modal integration to $P_{4}$ and $P_{5}$ while routing $P_{3}$ features exclusively from the RGB branch. Fig. 1 illustrates the complete topology.
@@
Figure 1: Overall architecture of CMAFNet. The dual-branch encoder processes RGB and depth modalities through symmetric but parameter-independent pathways, with SRM deployed at $P_{3}$ and $P_{4}$ for intra-modal feature purification. Cross-modal fusion occurs at $P_{4}$ and $P_{5}$; CSIF provides global attention-based context modeling at the $P_{5}$ fusion stage. The detection head implements bidirectional FPN-PAN feature propagation for multi-scale prediction. Depth features at $P_{3}$ are excluded from the detection pathway (dashed termination), so the highest-resolution prediction scale relies solely on RGB spatial detail.
The data flow comprises three stages. In the first, the dual-branch backbone encodes RGB and depth inputs independently through to $P_{5}$, with SRM inserted at $P_{3}$ and $P_{4}$ within each branch to suppress modality-specific noise and narrow the inter-modal distribution gap. In the second, cross-modal fusion at $P_{4}$ and $P_{5}$ combines the purified branch features through channel concatenation, C3k2 re-encoding, and a second SRM pass; at $P_{5}$, CSIF is additionally interposed between C3k2 and the terminal SRM to model global context. In the third, the detection head receives the fused $P_{4}$$P_{5}$ representations together with the RGB-only $P_{3}$ features through the standard FPN-PAN bidirectional pyramid for multi-scale prediction.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone comprises two encoding branches with identical topology but entirely separate parameters, one for RGB and one for depth. All convolutional layers, C3k2 blocks, and the terminal SPPF module adopt the YOLO 11(Jocher et al., 2023) default implementation without modification.
The RGB branch receives 3-channel images. The depth branch receives single-channel depth maps.Depth maps in the TL-RGBD dataset are spatially registered with their RGB counterparts via the factory calibration of the onboard RGB-D sensor (Intel RealSense D435i), and no additional geometric alignment is performed during preprocessing. Missing depth values and edge artifacts inherent to the sensor are not explicitly inpainted; robustness to these corruptions is instead delegated to SRM purification and the selective fusion strategy described in Section 2.1.2.
For the nano-scale configuration (width_multiple =0.25=0.25, depth_multiple =0.50=0.50, input $640\times 640$), the feature hierarchy spans five pyramid levels: $P_{1}$ ($320{\times}320$, C=16C{=}16), $P_{2}$ ($160{\times}160$, C=32C{=}32), $P_{3}$ ($80{\times}80$, C=64C{=}64), $P_{4}$ ($40{\times}40$, C=128C{=}128), and $P_{5}$ ($20{\times}20$, C=256C{=}256), with strides of 2, 4, 8, 16, and 32 respectively. The two branches produce identical feature dimensions at every level; the sole asymmetry is that the depth branch’s $P_{3}$ output is not propagated to the detection head.
Within each branch, an SRM instance (K=128K{=}128, $\alpha{=}0.8$; see Section 2.2) is inserted immediately after the C3k2 block at both $P_{3}$ and $P_{4}$. These scales strike a favorable balance: semantic content is rich enough for effective refinement, yet spatial resolution is still fine enough to preserve small-defect detail. Applying SRM before fusion narrows the inter-modal distribution gap at the source, providing cleaner inputs for the subsequent concatenation-based integration.
The rationale for parallel branches, rather than early fusion, lies in the distinct noise profiles of the two modalities. RGB imagery exhibits illumination variation and specular reflections; depth maps suffer from edge bleeding, missing values at reflective surfaces, and quantization noise. Merging these signals before individual processing would entangle their noise distributions and complicate subsequent alignment. Independent encoding through $P_{5}$ lets each branch develop modality-appropriate representations, deferring cross-modal integration to $P_{4}$$P_{5}$ where greater semantic abstraction makes noise-induced misalignment less problematic.
2.1.2 Fusion Layers: Selective Cross-Modal Integration at $P_{4}$ and $P_{5}$